A glass camera lens with an index of refraction of 1.55 is to be coated with a cryolite film (n = 1.3) to decrease the reflections of normally incident green light of wavelength 500 nm. How thick should such a film be?
Answer #1

Given:
Index of refraction of thin film of cryolite nc =1.3

Wavelength of light λ = 500*10-9m

Here we have the condition for brightness in case of reflection as
2nt = mλ. . . . .so set it equal to t
t = mλ / 2n
for minimum thickness m = 1
t = 1*500*10-9m / (2*1.3)
= 220.2nm

answered by: ANURANJAN SARSAM